mJy1,2019, the balance in the Supplies account was $400. During the year, $11,500 of supplies was purchased on account. 0n December 31, 2019 the company revealed that there were supplies of $1,100 remaining in the closet Required: a) What amount of Supplies Expense will be reported on the 2019 income statement? b) What account and amount, if any, will be reported on the 2019 Balance Sheet? Account Name 47. (Pains) This company paid $24.000 for sis monthsreat Angust 1. 2019. a) What amount if any, of Rent Expense will be shown on the 2019 income statement? b) What amount, if any, would be shonn on the December 31, 2019 balance sheet (do not inchude cash)? Account Name 48. ( Painrs The company issued a $100,000 6 month 5% promisory note on October 1,20 1g to Natan I Bank with interest paid at the end of each mont. a) Whar accounte) and amounte) wl be showa on the 2019 incone statem.ext Account Name Account Name b) Wbat accountrts), it any, would be shonn on the December 31. 2019 balance sheet? Do rot include cash Account Name Account Name
Answer #1

(a): Supplies expense = opening balance in supplies account+purchases – closing balance in supplies account

= 400+11500-1100

= $10,800

(b): Account Name: Supplies Account. $1,100

(47) a: Monthly rent = $24,000/6 months = $4,000 per month. So rent expenses from August to December = $4,000*5 months = $20,000

Thus rent expenses in 2019 income statement = $20,000

(b): Account name: Pre-paid rent. $4,000

(48) Monthly interest payments = $100,000*5%/12 = $416.67.

Thus interest expenses for October to December 2019 = $416.67*3 months = $1,250. This amount will not stand as interest payable as the amount is paid at the end of each month.

Thus accounts in income statement and the amount will be:

Account name: Interest expenses. Amount $1,250

Accounts and amounts in Balance sheet will be:

Accounts name: Notes payable. Amount $100,000

(Please note that as interest is paid at the end of each month hence it can be assumed that there are no interest payable and all interest amount is paid at the end of each month to the bank. Hence the balance sheet will not have an interest payable account).